You are dealt two cards from a deck of 52 cards. What is the probability that your two dealt cards are the 3 of clubs and the queen of hearts?
Answer #1

There is only one way of picking 3 of clubs & queen of hearts. Total number of ways of choosing 2 cards out of 52 distinct cards = (52C2) = (52*51)/2 = 1326

Hence required probability = 1/1326 = 0.00075415